# Testing disformal non-circular deformation of Kerr black holes with LISA

### v1

@@ I Introduction
EMRIs constitute a small compact object (SCO), (e.g. a stellar-mass black hole or a neutron star) spiraling into a supermassive black hole (SMBH) in a galactic center with masses of the order of $10^{5}-10^{7}M_{\odot}$. The large difference of mass scales, typically $10^{-6}-10^{-4}$, means that the SCO influence on the binary system may safely be considered to be a perturbation of the SMBH spacetime. The SCO is expected to undergo typically about $10^{5}-10^{6}$ cycles around the SMBH in the LISA frequency band before plunging. This fact is important for fundamental physics because even very small deviations caused by possibly new physics accumulate and eventually could be detected by LISA and other future space-based detectors. For example, EMRIs can probe interesting features such as the presence of a scalar field around the primary black hole [5, 6, 7] or even the secondary one [8], nonintegrable deformations [9], parity violation [10, 11, 12], etc. [13].
There is strong observational evidence that almost every large galaxy has a SMBH at its center. For example, our galaxy hosts a SMBH corresponding to the radio source Sagittarius A*. A fundamental question that needs to be answered is whether the central objects are described by the standard Kerr solution in General Relativity (GR) or by another black hole solution. Even though until now some more exotic geometries were ruled out mainly by the Event Horizon observations [14, 15], their accuracy is still not good enough to have a conclusive answer. The hope is that LISA will be able to fill this gap.
@@
The disformal Kerr metric is one of the very few stationary solutions that provides a measurable departure from Kerr. Being an explicit and simple enough solution, its properties can be studied in parallel with Kerr in a more or less straightforward fashion. Like all explicit solutions it has a number of positive but also negative points which stem from its explicit form and construction properties. It is constructed from the stealth Kerr solution [24] based on a scalar field which is a regular Hamilton-Jacobi functional depicting integrable Kerr geodesics. As such in stealth Kerr the scalar hair is very particular and paints the spacetime in a geodesic fashion. The tensor perturbations obey a modified Teukolsky equation [25] but the scalar perturbations are frozen and the scalar does not obey a hyperbolic type of equation [26, 27]. This most probably renders the scalar perturbations of the disformal Kerr solution pathological. On the positive side the metric is surprisingly regular, does not contain closed causal curves, and has other non trivial properties. This is thanks to the fact that the deformation is associated to geodesics of the Kerr metric itself [24]. Given the difficulty of obtaining explicit stationary spacetimes, ad hoc metrics are often discussed in the literature (see for example [28] and more recently [29] and references within). However, ad hoc metrics often do not share regularity and/or causality of the dismorfed Kerr metric being a consequence of the nature of their ad hoc construction.
@@ III.2 Gravitational waves
$A_{\times}=\frac{4\mu M}{{\cal D}_{L}}\Omega^{2}(t)R^{2}(t)\cos\Theta\sin\left(\Phi_{GW}(t)\right),$
@@ V Detecting small disformal deformations of Kerr black holes with LISA
The LISA sensitivity to the phase resolution of EMRI measurements is estimated as $\Delta\Phi_{GW}\sim 0.1$ assuming the signal to noise ratio (SNR) to be $SNR\geq 30$ [36]. Therefore the minimal disformal parameter $|D_{min}|$ that could in principle be detected by LISA is

### v2

@@ I Introduction
EMRIs constitute a small compact object (SCO), (e.g. a stellar-mass black hole or a neutron star) spiraling into a supermassive black hole (SMBH) in a galactic center with masses of the order of $10^{5}-10^{7}M_{\odot}$. The large difference of mass scales, typically $10^{-6}-10^{-4}$, means that the SCO influence on the binary system may safely be considered to be a perturbation of the SMBH spacetime. The SCO is expected to undergo typically about $10^{5}-10^{6}$ cycles around the SMBH in the LISA frequency band before plunging. This fact is important for fundamental physics because even very small deviations caused by possibly new physics accumulate and eventually could be detected by LISA and other future space-based detectors. For example, EMRIs can probe interesting features such as the presence of a scalar field around the primary black hole [5, 6, 7] or even the secondary one [8], the presence of dark matter [9], nonintegrable deformations [10, 11], parity violation [12, 13, 14], etc. [15, 16, 17].
There is strong observational evidence that almost every large galaxy has a SMBH at its center. For example, our galaxy hosts a SMBH corresponding to the radio source Sagittarius A*. A fundamental question that needs to be answered is whether the central objects are described by the standard Kerr solution in General Relativity (GR) or by another black hole solution. Even though until now some more exotic geometries were ruled out mainly by the Event Horizon observations [18, 19], their accuracy is still not good enough to have a conclusive answer. LISA will be able to fill this gap [3, 4].
@@
The disformal Kerr metric is one of the very few stationary solutions that provides a measurable departure from Kerr. Being an explicit and simple enough solution, its properties can be studied in parallel with Kerr in a more or less straightforward fashion. Like all explicit solutions it has a number of positive but also negative points which stem from its explicit form and construction properties. Disformal Kerr is constructed from the stealth Kerr solution [28] which is Kerr spacetime accompanied by a non trivial scalar field with a radial profile and a linear time dependence. It turns out [28], that the scalar field is associated to a geodesic congruence of Kerr. The scalar is in fact a particular Hamilton-Jacobi functional depicting integrable Kerr geodesics which is chosen to be regular. As such in stealth Kerr the scalar hair is very particular and accommodates spacetime in a geodesic and regular fashion. The tensor perturbations obey a modified Teukolsky equation [29] but the scalar perturbations are frozen and the scalar does not obey a hyperbolic type of equation [30, 31]. This most probably renders the scalar perturbations of the disformal Kerr solution pathological. On the positive side the metric is surprisingly regular, does not contain closed causal curves, and has other non trivial properties. This is thanks to the fact that the deformation is associated to geodesics of the Kerr metric itself [28]. Given the difficulty of obtaining explicit stationary spacetimes, ad hoc metrics are often discussed in the literature (see for example [32] and more recently [33] and references within). However, ad hoc metrics often do not share regularity and/or causality of the dismorfed Kerr metric being a consequence of the nature of their ad hoc construction.
@@ III.2 Gravitational waves
$A_{\times}=\frac{4\mu M}{{\cal D}_{L}}\Omega^{2}(t)R^{2}(t)\cos\Theta\sin\left(\Phi_{GW}(t)\right),$ (30)
@@ V Detecting small disformal deformations of Kerr black holes with LISA
The LISA sensitivity to the phase resolution of EMRI measurements is estimated as $\Delta\Phi_{GW}\sim 0.1\,{\rm rad}$ assuming the signal to noise ratio (SNR) to be $SNR\geq 30$ [40]. Therefore the minimal disformal parameter $|D_{min}|$ that could in principle be detected by LISA is